# Shifted multiplicative subgroups are not ratio sets

### v1

@@ Abstract.
In a recent breakthrough, Kalmynin proved a conjecture of Lev–Sonn and a conjecture of Sárközy on additive decompositions of multiplicative subgroups of a prime field. In this paper, we prove a multiplicative analogue of Kalmynin’s result on a generalization of the Lev–Sonn conjecture, inspired by a relevant conjecture of Sárközy. We show that all nonzero shifts of proper multiplicative subgroups (of size at least $3$) are not ratio sets of the form $A/A$. This in particular extends a result of Shkredov, where he showed the same for small multiplicative subgroups (of size $p^{6/7}$ in $\mathbb{F}_{p}$). We also prove an analogous statement over complex numbers for finite subgroups of the unit circle, which may be of independent interest.
@@
multiplicative subgroup, multiplicative decomposition
@@ 1. Introduction
In this paper, we study finite field analogues of these questions. In particular, we study certain multiplicative decompositions of shifted multiplicative subgroups of finite fields. The following well-known conjecture is due to Sárközy [12].
@@
It is natural to consider the analogue of this conjecture over multiplicative subgroups of finite fields. In particular, the following generalization of Sárközy’s conjecture is well-believed:
@@
Together with Ford’s theorem [4] on the distribution of divisors of shifted primes, Hanson and Petridis showed that Theorem 1.3 implies an asymptotic version of 1.1 (see [6, Corollary 1.4]). In a very recent breakthrough, using a sophisticated application of Stepanov’s method, together with various tools from number theory and combinatorics, Kalmynin [7] resolved Conjecture 1.1 and made substantial progress towards 1.2 over prime fields.
@@
In this paper, we study a multiplicative analogue of the difference set variant of Conjecture 1.1 proposed by Lev and Sonn [9]. We begin by recalling the Lev–Sonn conjecture.
@@
In [7], Kalmynin used an ingenious idea to confirm the generalization of 1.4 for multiplicative subgroups of prime fields.
@@
In fact, it is implicit in his proof [7, Section 3] that a slightly stronger statement holds: under the same assumption of Theorem 1.5, if $A-A\subseteq G\cup\{0\}$, then
@@
Next, we turn to multiplicative analogues of the conjectures and results discussed above, before discussing our main result. These analogues were also initiated by Sárközy [13].
@@
In [8], inspired by 1.2, we formulated a generalization of this multiplicative Sárközy conjecture for proper multiplicative subgroups.
Conjecture 1.7 ([8, Conjecture 1.9]).
@@
Using Stepanov’s method, we made some partial progress on 1.7 in our previous paper [8, Theorem 1.11]. The following lemma is a simplified version of [8, Theorem 1.1], which is the key ingredient in the proof of [8, Theorem 1.11] in the same paper.
Lemma 1.8 ([8, Theorem 1.1]).
@@
In particular, for $\lambda=1$, we confirmed 1.6 for almost all primes $p$ [8, Theorem 1.11]. We remark that the multiplicative Sárközy conjecture is potentially more difficult than the additive Sárközy conjecture  [^1] Private communication with Ilya Shkredov.. Indeed, in Shkredov’s paper [17], he made important progress on 1.2 by showing that all small multiplicative subgroups of prime fields have no nontrivial additive decomposition. By contrast, he obtained only a much weaker statement toward multiplicative decompositions of shifted multiplicative subgroups: if $\epsilon>0$ is fixed, then for all small multiplicative subgroups $G$ of ${\mathbb{F}}_{p}$ satisfying $1\ll_{\epsilon}|G|\ll_{\epsilon}p^{6/7-\epsilon}$,
@@
Similarly, when $\lambda\notin G$, the conclusion of Lemma 1.8 is weaker and not strong enough for our intended application to 1.7.
Note that Shkredov’s result confirms a multiplicative analogue of the generalized Lev–Sonn conjecture for small multiplicative subgroups. Our main result, stated below, establishes a multiplicative analogue of Kalmynin’s theorem (Theorem 1.5), and in particular extends the above result of Shkredov [17] to all proper multiplicative subgroups of order at least $3$. Our proof is based on a combination of Stepanov’s method and tools from symmetric polynomials. It is inspired by Kalmynin’s paper [7], although the exact proof techniques and the choice of auxiliary polynomials are different.
@@
    $A/A\neq\big((\xi G\cup\{0\})+\mu\big)\setminus\{0\}.$ (1.3)
We believe that both parts of the theorem are natural multiplicative analogues of Kalmynin’s theorem (Theorem 1.5). Also, we remark that the assumptions that $G$ is proper and $|G|\geq 3$ are necessary (in both parts of the above theorem) by considering the following counterexamples:
@@
  • $|G|=2$: we have $G=\{1,-1\}$ and we can take $A=\{1\}$ so that $A/A=(\frac{1}{2}G+\frac{1}{2})\setminus\{0\}$. By taking $A=\{1,-2\}$, we have $A/A=\{1,-1/2,-2\}=(-\frac{3}{2}G\cup\{0\}-\frac{1}{2})\setminus\{0\}$.
  • $G={\mathbb{F}}_{p}^{*}$: let $g$ be a primitive root in ${\mathbb{F}}_{p}$ and set $A=\{g^{j}:0\leq j\leq\frac{p-3}{2}\}$, then
    By taking $A={\mathbb{F}}_{p}^{*}$, we simply have $A/A=({\mathbb{F}}_{p}+1)\setminus\{0\}$.
Finally, let us consider the setting of roots of unity in the field of complex numbers. Let $\mathbb{C}$} be the field of complex numbers and ${\mathbb{C}}^{*}={\mathbb{C}}\setminus\{0\}$. Let $S^{1}$ denote the unit circle, that is, $S^{1}=\{z\in{\mathbb{C}}:|z|=1\}$. In Kalmynin’s paper [7], as well as the paper of Hanson and Petridis [6], they also considered analogues of Theorem 1.3 and Theorem 1.5 for finite subgroups of $S^{1}$. In a similar spirit, in the following theorem, we establish an analogue of Theorem 1.9 for finite subgroups of $S^{1}$.
@@
    For every $\xi,\mu\in{\mathbb{C}}^{*}$ and every $A\subseteq{\mathbb{C}}^{*}$,
@@
    For every $\xi,\mu\in{\mathbb{C}}^{*}$ and every $A\subseteq{\mathbb{C}}^{*}$,
@@
Similar to the remark following Theorem 1.9, the assumption $|G|\geq 3$ is also necessary.
@@
In Subsection 2.1, we recall a standard result on symmetric polynomials. In Subsection 2.2, we revisit a multiplicative analogue of the Hanson-Petridis polynomial introduced in our previous paper [8], which plays a crucial role in the proof of Lemma 1.8. In Section 3, we prove Theorem 1.9, establishing multiplicative analogues of the generalized Lev-Sonn conjecture. Finally, in Section 4, we present a short proof of Theorem 1.10.
2. Preliminaries
@@ 2.1. Symmetric polynomials
In this subsection, we recall some standard families of symmetric polynomials and fix our notation. Throughout, $n\geq 1$ is an integer and $x_{1},\dots,x_{n}$ are indeterminates.
@@
$ke_{k}(x_{1},x_{2},\dots,x_{n})=\sum_{i=1}^{k}(-1)^{i-1}e_{k-i}(x_{1},x_{2},\dots,x_{n})p_{i}(x_{1},x_{2},\dots,x_{n}),$ (2.1)
@@
2.2. A multiplicative analogue of Hanson–Petridis polynomials
In this subsection, we revisit a multiplicative analogue of the Hanson–Petridis polynomial developed in our previous paper [8], and derive useful properties from it.
Following [8, Section 4.1], there exists a unique solution $c_{1},c_{2},\ldots,c_{n}\in{\mathbb{F}}_{p}$ of the following system of equations:
[…]
We considered the following auxiliary polynomial defined as:
$r(n-1)+(m-r)n=mn-r\leq\deg f\leq n-1+|G|.$ (2.4)
Thus, when equality holds in 2.4, $f$ must have degree n1+|G|=mnrn-1+|G|=mn-r and
where
$C=\sum_{i=1}^{n}c_{i}a_{i}^{n-1+|G|}\neq 0.$
3. Proof of Theorem 1.9
In this section, we prove Theorem 1.9.
@@ 2.2. A multiplicative analogue of Hanson–Petridis polynomials
and we follow the same notations as in Subsection 2.2 by viewing $B=1/(\xi A)$.
@@
Thus, we have $n^{2}-n+1=|G|$, and equality holds in 2.4. Thus, the auxiliary polynomial $f$ defined in equation 2.3 also satisfies equation 2.5. It follows that we have the following equality of polynomials:
@@
On the middle of equation 3.2, for each $1\leq k\leq n^{2}$, the coefficient of $x^{k}$ is
@@
By 3.1, we have $n\geq 3$. Let $a,b\in A\setminus\{1\}$ with $a\neq b$. Then we have
@@
(2) Suppose otherwise that $\mu=-1$. We have $A/A+1\subseteq\xi G\cup\{0\}$. By 3.1, we have $n\geq 3$. Without loss of generality, we may assume that $1\in A$. Then for each $a\in A\setminus\{-1\}$, we have
@@
and taking their ratio of these two elements in $\xi G$, $a$ lies in $G$. Thus, $A\setminus\{-1\}\subseteq G$.
@@
Let $r=|(\xi A)^{-1}\cap(-\lambda A^{-1})|$. Then, $r=|\{(a,b)\in A\times A\colon a/(b\xi)=-\lambda\}|$, so $\lambda$ has exactly $r$ representations in $A/(\xi A)$. Note that if $a,b\in A$ such that $a/(\xi b)=-\lambda$, then $b/(\xi a)=-1/(\lambda\xi^{2})$. Thus, $1/(\lambda\xi^{2})$ has exactly $r$ representations in $A/(\xi A)$ as well. Also, $1/\xi$ has exactly $n$ representations in $A/(\xi A)$. Note that $\lambda=\mu/\xi$ and $1/(\lambda\xi^{2})=1/(\mu\xi)$. We have $\mu\neq\pm 1$ by 3.2, thus $\lambda,-1/(\lambda\xi^{2}),1/\xi$ are pairwise distinct.
@@
We follow the same notations as in Subsection 2.2 by viewing $B=1/(\xi A)$.
@@
Since $n^{2}-n=|G|$, the equality holds in 2.4. Thus, the auxiliary polynomial $f$ defined in equation 2.3 also satisfies equation 2.5. It follows that we have the following equality of polynomials:
@@
Multiplying both hand-sides of equation 3.9 by $f(x)=f(0)+x^{n}R(x)$ and using the facts that $f^{\prime}(x)$ is divisible by $x^{n-1}$ and $f(0)\neq 0$, we obtain that for each $0\leq k\leq n-2$,

### v2

# Multiplicative irreducibility of shifted multiplicative subgroups

@@ Abstract.
In a recent breakthrough, Kalmynin resolved conjectures of Lev–Sonn and Sárközy on additive decompositions of multiplicative subgroups of prime fields. In this paper, inspired by a related conjecture of Sárközy, we prove multiplicative analogues of Kalmynin’s results. We show that for every proper multiplicative subgroup $G$, the shifted set $(G-1)\setminus\{0\}$ cannot be written as a product set nontrivially, addressing a conjecture of Sárközy. In addition, we prove that no nonzero shift of any coset of a proper multiplicative subgroup is a ratio set of the form $A/A$. Our results substantially sharpen previous theorems of Shkredov and the authors.
@@
multiplicative subgroup, multiplicative decomposition, Stepanov’s method
@@ 1. Introduction
In this paper, we study finite field analogues of these questions. In particular, we study certain multiplicative decompositions of shifted multiplicative subgroups of finite fields. To motivate these questions and our main results, we begin with additive versions of these questions. The following well-known conjecture is due to Sárközy [13].
@@
It is natural to consider the analogue of this conjecture over multiplicative subgroups of finite fields. In particular, the following generalization of Sárközy’s conjecture is widely believed:
@@
Together with Ford’s theorem [4] on the distribution of divisors of shifted primes, Hanson and Petridis showed that Theorem 1.3 implies an asymptotic version of 1.1 (see [6, Corollary 1.4]).
In a very recent breakthrough, Kalmynin [8] used an ingenious idea to resolve Conjecture 1.1 and made substantial progress towards 1.2 over prime fields. More precisely, he proved the following theorem:
Theorem 1.4 (Kalmynin [8]).
Let $p$ be a prime and $G$ be a proper multiplicative subgroup of ${\mathbb{F}}_{p}$. If $A,B$ are subsets of ${\mathbb{F}}_{p}$ such that $A+B=G$ and $|A|,|B|\geq 2$, then $|G|$ is a perfect square and $|A|=|B|=\sqrt{|G|}$. Moreover, when $G=\mathcal{R}_{p}$, such sets $A,B$ do not exist.
To prove Theorem 1.4, he employed a sophisticated application of Stepanov’s method, building on the auxiliary polynomial introduced by Hanson and Petridis [6]. He also noted, in a talk at the Number Theory Web Seminar, that the techniques used to resolve 1.1 do not appear to extend to general multiplicative subgroups.
@@
In the same paper, using similar techniques, Kalmynin [8] resolved the following difference set variant of Conjecture 1.1 proposed by Lev and Sonn [10].
@@
Kalmynin [8] proved the following generalization of 1.5 for multiplicative subgroups of prime fields. We note that the case $|G|\leq p^{4/5-o(1)}$ in the following theorem has been proved by Shkredov [17].
@@
In fact, it is implicit in his proof [8, Section 3] that a slightly stronger statement holds: under the same assumption of Theorem 1.6, if $A-A\subseteq G\cup\{0\}$, then
@@
Next, we turn to multiplicative analogues of the conjectures and results discussed above, before discussing our contributions. These analogues were also initiated by Sárközy [14].
@@
In [9], inspired by 1.2, we formulated a generalization of 1.7 for proper multiplicative subgroups.
Conjecture 1.8 ([9, Conjecture 1.9]).
@@
Using Stepanov’s method, we made some partial progress on 1.8 in our previous paper [9, Theorem 1.11]. The following lemma is a simplified version of [9, Theorem 1.1], which is the key ingredient in the proof of [9, Theorem 1.11] in the same paper.
Lemma 1.9 ([9, Theorem 1.1]).
@@
In particular, for $\lambda=1$, we showed that Lemma 1.9 implies an asymptotic version of 1.7 in [9, Theorem 1.11]. We also showed that it implies weaker versions of 1.8 in [9, Corollary 6.4 and Theorem 6.6]: if $G$ is a proper multiplicative subgroup of a prime field ${\mathbb{F}}_{p}$ with $|G|$ sufficiently large, then for all $\lambda\in{\mathbb{F}}_{p}^{*}$, the set $(G-\lambda)\setminus\{0\}$ cannot be written as a product set $AA$ or $ABC$ nontrivially.
We remark that the multiplicative Sárközy conjecture is potentially more difficult than the additive Sárközy conjecture  [^1] Private communication with Ilya Shkredov.. Indeed, in Shkredov’s paper [18], he made important progress on 1.2 by showing that all small multiplicative subgroups of prime fields have no nontrivial additive decomposition. By contrast, he obtained only a much weaker statement toward multiplicative decompositions of shifted multiplicative subgroups: if $\epsilon>0$ is fixed, then for all small multiplicative subgroups $G$ of ${\mathbb{F}}_{p}$ satisfying $1\ll_{\epsilon}|G|\ll_{\epsilon}p^{6/7-\epsilon}$,
@@
Similarly, when $\lambda\notin G$, the conclusion of Lemma 1.9 is weaker and not strong enough for our intended application to 1.8.
Our first main result resolves 1.8 for prime fields in the setting $\lambda\in G$.
Theorem 1.10.
Let $p$ be an odd prime, $G$ be a proper subgroup of ${\mathbb{F}}_{p}^{*}$, and $\lambda\in G$. Then there do not exist sets $A,B\subseteq{\mathbb{F}}_{p}^{*}$ with $|A|,|B|\geq 2$ such that
$AB=(G-\lambda)\setminus\{0\}.$
Theorem 1.10 can be viewed as a multiplicative analogue of Theorem 1.4: it is somewhat stronger in the sense that Theorem 1.4 does not resolve the general subgroup setting; however, it is somewhat weaker in the sense that we need the additional assumption $\lambda\in G$. We believe that the case $\lambda\notin G$ is substantially more difficult, and our techniques do not seem to extend to the case $\lambda\in{\mathbb{F}}_{p}^{*}\setminus G$. We also remark that Theorem 1.10 does not always hold when $\lambda\in{\mathbb{F}}_{p}^{*}\setminus G$. Indeed, a short computer search finds the following two counterexamples:
  1. (1)
    In ${\mathbb{F}}_{11}$, we have $(G-\lambda)\setminus\{0\}=AB$ for
    $G=\{1,3,4,5,9\},\quad\lambda=2,\quad A=\{1,7\},\quad B=\{1,2,3\}.$
  2. (2)
    In ${\mathbb{F}}_{19}$, we have $(G-\lambda)\setminus\{0\}=AB$ for
    $G=\{1,7,8,11,12,18\},\quad\lambda=2,\quad A=\{1,9\},\quad B=\{6,9,18\}.$
We also note that the assumption that $G$ is proper in Theorem 1.10 is necessary. Indeed, let $g$ be a primitive root in ${\mathbb{F}}_{p}$ and set $A=\{g^{j}:0\leq j\leq\frac{p-3}{2}\}$; then
Next, we consider the simpler setting of expressing a shifted multiplicative subgroup as a ratio set $A/A$. Recall that Shkredov’s result confirms a multiplicative analogue of the generalized Lev–Sonn conjecture for small multiplicative subgroups. Our second main result, stated below, completely establishes a multiplicative analogue of Theorem 1.6, and in particular extends the above result of Shkredov [18] to all proper multiplicative subgroups of order at least $3$.
@@
    $A/A\neq\big((\xi G\cup\{0\})+\mu\big)\setminus\{0\}.$
We believe that both parts of the theorem are natural multiplicative analogues of Theorem 1.6. Also, we remark that the assumptions that $G$ is proper and $|G|\geq 3$ are necessary (in both parts of the above theorem) by considering the following counterexamples:
@@
  • $|G|=2$: we have $G=\{1,-1\}$ and we can take $A=\{1\}$ so that $A/A=(\frac{1}{2}G+\frac{1}{2})\setminus\{0\}$. By taking $A=\{1,-2\}$, we have $A/A=\{1,-1/2,-2\}=((-\frac{3}{2}G\cup\{0\})-\frac{1}{2})\setminus\{0\}$.
  • $G={\mathbb{F}}_{p}^{*}$: the remark following Theorem 1.10 shows it is possible to have $A/A=({\mathbb{F}}_{p}^{*}-1)\setminus\{0\}.$ By taking $A={\mathbb{F}}_{p}^{*}$, we simply have $A/A=({\mathbb{F}}_{p}+1)\setminus\{0\}$.
Our proofs of Theorems 1.10 and 1.11 rely on a combination of Stepanov’s method and tools from symmetric polynomials. While they are motivated by Kalmynin’s paper [8], the actual arguments are quite different, and the resulting proofs seem to be simpler. We also note that our methods apply more generally to small multiplicative subgroups of arbitrary finite fields. For example, with only minor modifications, the proofs of Theorems 1.10 and 1.11 extend to multiplicative subgroups $G$ of finite fields ${\mathbb{F}}_{q}$ satisfying $|G|\leq p-2\sqrt{p}$, where $p$ is the characteristic of ${\mathbb{F}}_{q}$.
Finally, we consider the setting of roots of unity in the field of complex numbers. Let $\mathbb{C}$} be the field of complex numbers and ${\mathbb{C}}^{*}={\mathbb{C}}\setminus\{0\}$. Let $S^{1}$ denote the unit circle, that is, $S^{1}=\{z\in{\mathbb{C}}:|z|=1\}$. In Kalmynin’s paper [8], as well as the paper of Hanson and Petridis [6], they also considered analogues of Theorems 1.3 and 1.6 for finite subgroups of $S^{1}$. In a similar spirit, in the following theorem, we establish an analogue of Theorems 1.10 and 1.11 for finite subgroups of $S^{1}$. Note that in this case, the proof is much easier compared to the finite field setting, as we can apply a geometric argument.
@@
    Let $\lambda\in G$. Then there do not exist sets $A,B\subseteq{\mathbb{C}}^{*}$ with $|A|,|B|\geq 2$ such that
    $AB=(G-\lambda)\setminus\{0\}.$
  2. (2)
    If $|G|\geq 3$, then for every $\xi,\mu\in{\mathbb{C}}^{*}$ and every $A\subseteq{\mathbb{C}}^{*}$,
@@
    If $|G|\geq 3$, then for every $\xi,\mu\in{\mathbb{C}}^{*}$ and every $A\subseteq{\mathbb{C}}^{*}$,
@@
Similar to the remark following Theorem 1.11, the assumption $|G|\geq 3$ is also necessary in (2) and (3).
@@
In Section 2, we revisit a multiplicative analogue of the Hanson–Petridis polynomial introduced in our previous paper [9], which plays a crucial role in the proof of Lemma 1.9. In Section 3, we prove Theorem 1.11, establishing multiplicative analogues of the generalized Lev–Sonn conjecture. In Section 4, we resolve multiplicative analogues of the generalized Sárközy conjecture in the case $\lambda\in G$ (Theorem 1.10). Finally, in Section 5, we present a short proof of Theorem 1.12.
2. A multiplicative analogue of Hanson–Petridis polynomials
In this section, we revisit a multiplicative analogue of the Hanson–Petridis polynomial developed in our previous paper [9], and derive useful properties from it.
Following [9, Section 4.1], there exists a unique solution $c_{1},c_{2},\ldots,c_{n}\in{\mathbb{F}}_{p}$ to the following system of equations:
[…]
The following lemma will be a key observation in the proof of Theorem 1.10.
Lemma 2.1.
The coefficients $c_{1},c_{2},\dots,c_{n}$ satisfy
$\sum_{i=1}^{n}\frac{c_{i}a_{i}^{n}}{1-a_{i}x}=(-1)^{n-1}\Big(\prod_{i=1}^{n}a_{i}\Big)\cdot\frac{1}{\prod_{i=1}^{n}(1-a_{i}x)}$
as an identity of rational functions in ${\mathbb{F}}_{p}(x)$.
Proof.
First, we compute the coefficients $c_{1},\dots,c_{n}$ explicitly in terms of the Vandermonde matrix. To do so, define
$W=\begin{bmatrix}1&a_{1}&a_{1}^{2}&\cdots&a_{1}^{n-1}\\ 1&a_{2}&a_{2}^{2}&\cdots&a_{2}^{n-1}\\ \vdots&\vdots&\ddots&\vdots\\ 1&a_{n}&a_{n}^{2}&\cdots&a_{n}^{n-1}\\ \end{bmatrix}$.
Then, $W$ is invertible and $W^{t}[c_{1},c_{2},\ldots,c_{n}]^{t}=[1,0,\ldots,0]^{t}$, so
$c_{1},\ldots,c_{n}]^{t}=(W^{-1})^{t}[1,0,\ldots,0]^{t},$
which is precisely the first row of $W^{-1}$. Therefore, using the formula for the inverse of a Vandermonde matrix (see for example [7, Section 0.9.11]), we obtain
$c_{i}=\frac{(-1)^{n-1}a_{1}a_{2}\cdots a_{n}}{a_{i}\cdot\prod_{j\neq i}(a_{i}-a_{j})}.$ (2.2)
Using equation 2.2, we have
$\sum_{i=1}^{n}\frac{c_{i}a_{i}^{n}}{1-a_{i}x}$ $=\sum_{i=1}^{n}\frac{(-1)^{n-1}\prod_{t=1}^{n}a_{t}}{a_{i}\prod_{j\neq i}(a_{i}-a_{j})}\cdot\frac{a_{i}^{n}}{1-a_{i}x}$
$=(-1)^{n-1}\Big(\prod_{t=1}^{n}a_{t}\Big)\sum_{i=1}^{n}\frac{a_{i}^{n-1}}{\prod_{j\neq i}(a_{i}-a_{j})}\cdot\frac{1}{1-a_{i}x}.$ (2.3)
Now, write
$\frac{1}{\prod_{j=1}^{n}(1-a_{j}x)}=\frac{1}{\prod_{j=1}^{n}a_{j}}\cdot\frac{1}{\prod_{j=1}^{n}(a_{j}^{-1}-x)}.$
By the partial fraction decomposition, we have
$\frac{1}{\prod_{j=1}^{n}(a_{j}^{-1}-x)}=\sum_{i=1}^{n}\frac{1}{a_{i}^{-1}-x}\cdot\frac{1}{\prod_{j\neq i}(a_{j}^{-1}-a_{i}^{-1})}.$ (2.4)
Note that for a fixed $i$, we have
$\prod_{j\neq i}(a_{j}^{-1}-a_{i}^{-1})=\prod_{j\neq i}\frac{a_{i}-a_{j}}{a_{i}a_{j}}=\frac{\prod_{j\neq i}(a_{i}-a_{j})}{a_{i}^{n-1}\prod_{j\neq i}a_{j}},$
thus
$\frac{1}{\prod_{j\neq i}(a_{j}^{-1}-a_{i}^{-1})}=\frac{a_{i}^{n-1}\prod_{j\neq i}a_{j}}{\prod_{j\neq i}(a_{i}-a_{j})}.$
Substituting these identities into equation (2.4) yields
$\frac{1}{\prod_{j=1}^{n}(a_{j}^{-1}-x)}=\Big(\prod_{j=1}^{n}a_{j}\Big)\sum_{i=1}^{n}\frac{a_{i}^{n-1}}{\prod_{j\neq i}(a_{i}-a_{j})}\cdot\frac{1}{1-a_{i}x}.$
Multiplying both sides by $\big(\prod_{j=1}^{n}a_{j}\big)^{-1}$ gives
$\frac{1}{\prod_{j=1}^{n}(1-a_{j}x)}=\sum_{i=1}^{n}\frac{a_{i}^{n-1}}{\prod_{j\neq i}(a_{i}-a_{j})}\cdot\frac{1}{1-a_{i}x}.$
Thus, comparing with equation 2.3, the lemma follows. ∎
Following [9, Section 4.1], consider the following auxiliary polynomial
$r(n-1)+(m-r)n=mn-r\leq\deg f\leq n-1+|G|.$ (2.6)
Thus, when equality holds in 2.6, $f$ must have degree n1+|G|=mnrn-1+|G|=mn-r and
where
$C=\sum_{i=1}^{n}c_{i}a_{i}^{n-1+|G|}\neq 0.$
Next, additionally assume that $AB+\lambda\subseteq G$ and $\lambda\in G$, which applies to the proof of Theorem 1.10 in Section 4. In this case, we have $r=0$. Indeed, if $b\in B\cap(-\lambda A^{-1})$, then we have $ab+\lambda=0$ for some $a\in A$, contradicting $0\notin G$. Thus, by the above discussion, each of $b_{1},\dots,b_{m}$ is a root of $f$ with multiplicity at least $n$. Also, observe that system 2.1 implies that
$f(0)=-\lambda^{n-1}+\sum_{i=1}^{n}c_{i}\lambda^{n-1+|G|}=-\lambda^{n-1}+\lambda^{n-1}=0,$
as well as that the coefficient of $x^{j}$ of $f$ is 0 for $1\leq j\leq n-1$. Thus, 0 is also a root of $f$ with multiplicity at least $n$. By comparing the sum of multiplicities of the roots $0,b_{1},b_{2},\dots,b_{m}$ of $f$ and the degree of $f$, we have the following stronger inequality:
$(m+1)n\leq|G|+n-1.$ (2.8)
Moreover, when equality holds in 2.8, mn=|G|1mn=|G|-1 and we have the following factorization of $f$:
$f(x)=C\cdot x^{n}\prod_{j=1}^{m}(x-b_{j})^{n}=C\Bigl(x\prod_{j=1}^{m}(x-b_{j})\Bigr)^{n}.$ (2.9)
where
$C=\sum_{i=1}^{n}c_{i}a_{i}^{n-1+|G|}\neq 0.$
3. Proof of Theorem 1.11
In this section, we prove Theorem 1.11.
@@ 3.1. Symmetric polynomials
We first recall some standard families of symmetric polynomials and fix our notation. Throughout, $n\geq 1$ is an integer and $x_{1},\dots,x_{n}$ are indeterminates.
@@
$ke_{k}(x_{1},x_{2},\dots,x_{n})=\sum_{i=1}^{k}(-1)^{i-1}e_{k-i}(x_{1},x_{2},\dots,x_{n})p_{i}(x_{1},x_{2},\dots,x_{n}),$
@@
and we follow the same notations as in Section 2 by viewing $B=1/(\xi A)$.
@@
Thus, we have $n^{2}-n+1=|G|$, and equality holds in 2.6. Thus, the auxiliary polynomial $f$ defined in equation 2.5 also satisfies equation 2.7. It follows that we have the following equality of polynomials:
@@
In the middle expression of equation 3.2, for each $1\leq k\leq n^{2}$, the coefficient of $x^{k}$ is
@@
By 3.2, we have $n\geq 3$. Let $a,b\in A\setminus\{1\}$ with $a\neq b$. Then we have
@@
(2) Suppose otherwise that $\mu=-1$. We have $A/A+1\subseteq\xi G\cup\{0\}$. By 3.2, we have $n\geq 3$. Without loss of generality, we may assume that $1\in A$. Then for each $a\in A\setminus\{-1\}$, we have
@@
and taking the ratio of these two elements of $\xi G$, $a$ lies in $G$. Thus, $A\setminus\{-1\}\subseteq G$.
@@
Let $r=|(\xi A)^{-1}\cap(-\lambda A^{-1})|$. Then,
$r=|\{(a,b)\in A\times A\colon a/(b\xi)=-\lambda\}|$
and thus $\lambda$ has exactly $r$ representations in $A/(\xi A)$. Note that if $a,b\in A$ such that $a/(\xi b)=-\lambda$, then $b/(\xi a)=-1/(\lambda\xi^{2})$. Thus, $1/(\lambda\xi^{2})$ has exactly $r$ representations in $A/(\xi A)$ as well. Also, $1/\xi$ has exactly $n$ representations in $A/(\xi A)$. Note that $\lambda=\mu/\xi$ and $1/(\lambda\xi^{2})=1/(\mu\xi)$. We have $\mu\neq\pm 1$ by 3.3, thus $\lambda,-1/(\lambda\xi^{2}),1/\xi$ are pairwise distinct.
@@
We follow the same notations as in Section 2 by viewing $B=1/(\xi A)$.
@@
Since $n^{2}-n=|G|$, the equality holds in 2.6. Thus, the auxiliary polynomial $f$ defined in equation 2.5 also satisfies equation 2.7. It follows that we have the following equality of polynomials:
@@
Multiplying both sides of equation 3.9 by $f(x)=f(0)+x^{n}R(x)$ and using the facts that $f^{\prime}(x)$ is divisible by $x^{n-1}$ and $f(0)\neq 0$, we obtain that for each $0\leq k\leq n-2$,